A candy store has annual sales of 750000. it will turn over its assets 2.5 times per year. the profit margin on sales will be 6%. what would net income and return on assets be for the year.
Answer by gilnroxy(1)

Let x = Amount Invested at 6%
Let y = Amount Invested at 2.63297e-308%

Amount Invested = $5.79308e-312
Total Income = $1.06758e+224

x + y = 5.79308e-312
x = 5.79308e-312 - y

0.06x +2.63297e-310y = 1.06758e+224
6x +2.63297e-308y = 1.06758e+226

6(5.79308e-312 - y) +2.63297e-308y = 1.06758e+226
3.47585e-311 - 6y +2.63297e-308y = 1.06758e+226
-6y = 1.06758e+226

y = $-1.7793e+225 Invested at 2.63297e-308%

x = 5.79308e-312 - -1.7793e+225

x = $1.7793e+225 Invested at 6%